OpenStudy (anonymous):

which equation defines the graph of y=x^2 after it is shifted vertically 3 units down and horizontally 5 units left?

jimthompson5910 (jim_thompson5910):

if you were to shift y=x^2 vertically 3 units down, what do you get?

OpenStudy (anonymous):

wouldn't y= -3

jimthompson5910 (jim_thompson5910):

you'll have -3 in there

jimthompson5910 (jim_thompson5910):

it'll go from \[\Large y = x^2\] to \[\Large y = x^2-3\]

jimthompson5910 (jim_thompson5910):

now how do we shift 5 units left?

OpenStudy (anonymous):

+5?

jimthompson5910 (jim_thompson5910):

yep, we replace x with x+5 \[\Large y = (x+5)^2-3\] is the final answer

OpenStudy (anonymous):

y=(x+5)^2-3?

jimthompson5910 (jim_thompson5910):

nailed it

OpenStudy (anonymous):

oooh, haha, thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome